# Recognising, Anticipating, and Mitigating LLM Pollution of Online Behavioural Research

### v1

@@ Abstract
Online behavioural research faces an emerging threat as participants increasingly turn to large language models (LLMs) for advice, translation, or task delegation: LLM Pollution. We identify three interacting variants through which LLM Pollution threatens the validity and integrity of online behavioural research. First, Partial LLM Mediation occurs when participants make selective use of LLMs for specific aspects of a task, such as translation or wording support, leading researchers to (mis)interpret LLM-shaped outputs as human ones. Second, Full LLM Delegation arises when agentic LLMs complete studies with little to no human oversight, undermining the central premise of human-subject research at a more foundational level. Third, LLM Spillover signifies human participants altering their behaviour as they begin to anticipate LLM presence in online studies, even when none are involved. While Partial Mediation and Full Delegation form a continuum of increasing automation, LLM Spillover reflects second-order reactivity effects. Together, these variants interact and generate cascading distortions that compromise sample authenticity, introduce biases that are difficult to detect post hoc, and ultimately undermine the epistemic grounding of online research on human cognition and behaviour. Crucially, the threat of LLM Pollution is already co-evolving with advances in generative AI, creating an escalating methodological arms race. To address this, we propose a multi-layered response spanning researcher practices, platform accountability, and community efforts. As the challenge evolves, coordinated adaptation will be essential to safeguard methodological integrity and preserve the validity of online behavioural research.
Raluca Rilla1†, Tobias Werner1†, Hiromu Yakura1†, Iyad Rahwan1, Anne-Marie Nussberger1†∗
@@
Fast, scalable access to diverse participants through online recruitment platforms such as Prolific or MTurk has enabled a revolution in behavioural research (Gosling and Mason, 2015; Goodman and Wright, 2022; Palan and Schitter, 2018). The promise of “100% human” responses lies at the core of their appeal for researchers seeking to study authentic human cognition and behaviour. [^1] For instance, Prolific.com, one of the leading survey providers, advertises its services as “Easily examine human actions and behaviors with our 100% human, ID-checked participants” (see https://www.prolific.com/academic-researchers, last accessed July 13th, 2025. Yet as large language models (LLMs) become increasingly capable and readily available, this promise is under growing threat, evidenced by participants themselves reporting routinely using LLMs to assist with study tasks (Zhang et al., 2025). We term the emerging phenomenon in which LLMs are involved in online tasks intended to measure human responses LLM Pollution.
Recent experiences from our lab illustrate the immediacy of this threat: we repeatedly observed evidently LLM-mediated responses in up to 45% of submissions, characterised by overly verbose summaries of instructions or distinctly non-human phrases such as “I don’t experience confusion in the same way humans do” in open-ended survey questions. While threats to data quality from automated bots are not entirely new, earlier concerns focused on technically less sophisticated and more easily detectable bots (Goodrich et al., 2023; Griffin et al., 2022; Pozzar et al., 2020; Storozuk et al., 2020; Xu et al., 2022; Cresci, 2020). LLMs’ growing fluency and accessibility have, however, substantially amplified the problem by generating responses that are increasingly indistinguishable from those written by humans (Liem, 2025). As such, LLM Pollution can arise through a range of participant behaviours (Zhang et al., 2025). Some may use LLMs innocently, for example, to translate instructions or to improve fluency. Others might rely on them more strategically, using them to automate responses or reduce effort on tedious tasks. In more extreme cases, participants may delegate the completion of entire studies to agentic LLMs, such as OpenAI’s Operator, or open-source tools like Browser-use that can operate browsers, interpret screenshots, and navigate experiments without human oversight.
@@
Box 1 — A Case Study in LLM Pollution Detection and Incidence In a recent study we conducted with participants recruited via Prolific, we found that a lack of bot protection measures led to an alarming rate of seemingly bot-generated answers. In our initial pilot study, we only tracked copying and pasting on a page containing an open-ended question. Notably, 45% of participants engaged in one or both of these actions, suggesting potential cases of Partial LLM Mediation and Full LLM Delegation. Even to an untrained eye, some of these responses were obviously generated by LLMs, representing, for instance, cases of overly-abstract summarisations of all previous instructions. Given these results, we implemented the following measures for the main study: Implemented two reCAPTCHA v2 checkpoints at the beginning and middle of our survey, asking participants to confirm that they are not bots by checking a box and potentially completing a short validation test (0.2% failed to advance); Activated Qualtrics’ version of reCAPTCHA v3, which uses information about users’ behaviour and past activity to assign each a score between 0 and 1, with lower scores indicating likely bot activity (2.7% participants scored lower than 0.7, a threshold we determined on the basis of our piloting, with the lowest score being 0.2); Added a honeypot question that included white text invisible to humans, but readable to bots scraping all content (1.6% participants included our study-irrelevant keyword “hazelnut”; out of these 16 participants, only 2 had low reCAPTCHA v3 scores); Prevented copy-pasting, while still tracking which participants attempted these actions (4.7% participants attempted to copy and/or paste text). While these incidence rates were less concerning than those observed during piloting, we still identified identical nonsensical responses from a few different participants who otherwise appeared legitimate, having passed our honeypot question, having obtained perfect reCAPTCHA scores, and not having been flagged for copying/pasting. This suggests that even the above measures were insufficient for detecting different forms of LLM Pollution, further emphasising the scope of the problem.
In the following, we identify three key variants through which LLM Pollution manifests: Partial LLM Mediation, Full LLM Delegation, and LLM Spillover. We then outline a set of mitigation strategies, ranging from practical design choices to broader institutional safeguards. Finally, we offer an outlook on the future of online experiments in an era where human and machine responses are increasingly entangled.
@@
Figure 1: Three Variants of LLM Pollution. Partial LLM Mediation refers to cases where participants use LLMs for translation, idea generation, or performance gains. Full LLM Delegation involves the use of agentic tools or plugins to automate participation entirely. LLM Spillover refers to normative or behavioural shifts prompted by participants’ beliefs about LLM involvement, regardless of whether any is present. These variants can interact and reinforce one another, compounding their threat to research validity and inference.
@@
First, Partial LLM Mediation arises when participants rely on LLMs to help refine or generate their responses. Such responses may then appear to be human-generated but are, in part, machine-shaped. For instance, participants may rely on LLMs to translate content, improve the fluency of their writing, reduce cognitive effort, provide strategic advice on how to complete the task efficiently, or to maximise payment. Participants may copy and paste study instructions into LLM interfaces such as ChatGPT, or use screen-reading agents like Gemini Live to receive guidance and mediate their participation in the experiment. In some open-ended responses, LLM involvement may be evident from overly fluent language or stylistic quirks typical of LLMs (see, for instance, Yakura et al., 2024; Liang et al., 2024). Still, in many cases, Partial LLM Mediation remains difficult to detect from output alone, particularly when assistance is used only sparsely or intermittently. Partial LLM Mediation can compromise inference in several ways: because LLM-generated responses tend to exhibit lower variance and reflect dominant patterns in training data, their inclusion may artificially inflate central tendencies and obscure the true distribution of human responses (Zhang et al., 2025). These dominant patterns often encode systematic biases, such as the overrepresentation of Western, Educated, Industrialised, Rich, and Democratic (WEIRD) linguistic and cultural norms (Atari et al., 2023; Tao et al., 2024), which in turn are posed to produce homogenised outputs that mask meaningful variation, even within diverse participant samples. Most critically, Partial LLM Mediation undermines the validity of participant-generated responses. Researchers may believe they are analysing purely human-generated data, when in fact they are observing a blend of human and AI responses. This leads to incorrect inferences when the goal is to understand human cognition or behaviour, or to contrast it with AI behaviour.
@@
The second variant, Full LLM Delegation, involves participants outsourcing the entire study interaction to LLM-based tools or agents. These systems combine a language model with a control interface that enables goal-directed interaction with web environments. In practice, browser-integrated LLM agents (e.g., OpenAI’s Operator, open-source solutions like Browser-Use, Skyvern, or Nanobrowser) can autonomously read study content, interpret instructions, click through consent forms, and complete questionnaires, and generate both structured and open-ended responses — all with minimal or no human intervention once launched. Moreover, these agentic LLMs are increasingly capable of identifying instructions that are designed to mislead or manipulate their behaviour (see Figure B3). They are built to operate autonomously but are also designed to permit temporary human handover when they encounter barriers, such as CAPTCHA or other bot-detection systems. Furthermore, participants often have the option to instruct LLMs to adopt specific standpoints, personas, or tones that make their responses appear more credible both in terms of content and formatting at the surface level (Höhne et al., 2025a), further obscuring their underlying homogeneity (Zhang et al., 2025). Full LLM Delegation amplifies several risks observed in Partial LLM Mediation, as it more fully caters to participants’ motivations for efficiency and convenience, while distancing the resulting responses even further from genuine human behaviour. Hence, it first and foremost deepens the epistemic threat: researchers can no longer assume responses reflect human cognition, even probabilistically. Second, Full LLM Delegation may undermine treatment orthogonality, as LLMs may systematically respond differently across experimental conditions — especially in novel tasks — based on learned heuristics from training data (Mei et al., 2024; Schmidt et al., 2024). Third, Full LLM Delegation enables scale: participants (or third parties) can deploy LLM agents across multiple studies simultaneously, transforming data collection platforms into sites of automated labour arbitrage. As these systems continue to evolve and adapt rapidly, Full LLM Delegation is likely to become more common and harder to detect, posing a growing challenge for future research. Though described separately for clarity, Partial LLM Mediation and Full LLM Delegation lie on a continuum of increasing automation, with gradual shifts and no clear cutoffs between levels of LLM involvement.
The third variant of LLM Pollution focuses not on how participants use LLMs, but on how they respond to the perceived presence of LLMs in online research environments — what we term LLM Spillover. As LLM usage becomes more visible and suspected in online studies, participants may begin to change their behaviour simply in anticipation of LLM involvement, even if none is present. For example, many experiments include team tasks where participants interact with one another. If participants suspect they are paired with a bot rather than a human, this can influence their level of cooperation, trust, or engagement, regardless of whether an LLM is actually present. Some participants may try to signal that they are human by deliberately introducing typos or idiosyncratic phrasing to try to distinguish themselves from LLM-generated content. This type of signalling may not only target other participants but also the experimenter, introducing a new kind of demand effect where participants alter their behaviour based on what they believe the researcher expects or is trying to detect. Conversely, other participants may reduce their effort, rationalising that “everyone is cheating with LLMs anyway,” leading to moral licensing behaviours. In studies with sensitive questions, some participants might hold back their true opinions because they worry an LLM is observing or judging their answers. When LLM usage is perceived as widespread, some may justify their own use by appealing to a presumed norm. Even in the absence of actual machine involvement, these anticipatory behaviours can introduce noise, bias, and systematic distortion, posing a unique challenge to the interpretability of online research.
@@
Since efforts to detect LLM Pollution are already headed towards an arms race — as tools improve, so too do evasion strategies — eliminating LLM-generated responses entirely is unlikely to be feasible. Instead, we propose pragmatic, multi-layered mitigation strategies aimed at raising the cost and lowering the feasibility of LLM Pollution across different levels of behaviour research ecosystems. This includes safeguards that individual researchers can implement in study design, greater responsibility on research platforms to ensure sample integrity, and broader efforts to foster community-wide norms that promote data quality. Crucially, different forms of LLM Pollution require tailored responses, as no single tool addresses all threats equally (Höhne et al., 2025b).
@@
Measure Example Implementations Responsibility Purpose Variants Limitations
@@
Input interface restriction Block copying and pasting, or collecting responses via audio recording Researchers Prevention LLM Mediation LLM Delegation LLM Spillover Audio recording might have privacy concerns despite its effectiveness.
@@
Honeypot questions Add very small or white text with specific instructions for LLMs Researchers Post-hoc detection LLM Mediation LLM Delegation Advanced LLM-based agents might be able to detect and ignore honeypot questions.
@@
One set of strategies focuses on preventative measures designed to increase friction for LLM use. Commercial bot protection tools like reCAPTCHA or Cloudflare are widely available and relatively easy to implement. These tools can block access from known LLM platforms or agents (e.g., OpenAI’s Operator), and the platforms themselves are less likely to override such protections for ethical reasons. However, locally run tools or hybrid user-agent models can often bypass these barriers by temporarily returning control to a human. Another line of defence is to vary the mode of instruction presentation by using screenshots, images, or short videos to reduce the ease of copying and pasting into LLMs. While this may limit casual use, vision-enabled models and screen-sharing tools can again circumvent such barriers. Relatedly, restricting input methods can make automation more difficult, by means of disabling pasting, requiring typed responses, or blocking right-click functionality. Also, collecting responses via voice input may offer an additional layer of friction, as it is currently harder to automate reliably. Importantly, these strategies must be weighed against usability and privacy concerns, particularly in longer studies or with vulnerable populations. Finally, researchers may design tasks that contemporary LLMs struggle with, such as visual illusions or Theory of Mind questions, combined with additional comprehension checks (see Figures A1 and A2 for examples). While current models tend to overcorrect or produce obviously wrong answers in such tasks, these vulnerabilities are likely to diminish as models improve. As such, preventative strategies may offer only temporary protection unless regularly updated.
@@
Beyond prevention, researchers can also take steps to identify LLM involvement after the fact. Honeypot questions, such as inserting white or inaccessible text, can detect whether a participant’s browser is scraping the entire page, as many LLM agents do. These traps exploit the tendency of automated tools to process all on-screen content indiscriminately, including elements that human users cannot see (see Appendix A.7 for examples). Recording behavioural data like typing speed, mouse movement, and tab switching can also flag suspicious patterns. While helpful, these signals can be mimicked by more advanced agents. Researchers may choose to run AI-generated text detectors on open-ended responses. These tools, such as GPTZero or Undetectable AI, can provide estimates of LLM involvement, but they often lack transparency and may quickly become outdated as models evolve.
@@
Mitigating human reactivity effects that come to bear in LLM Spillover poses a distinct challenge, as interventions themselves can influence participant expectations. Simply informing participants that bot detection tools are in use may discourage casual or opportunistic reliance on LLMs, and explicit reminders that responses should reflect their own judgment can reduce mild or exploratory use. However, these interventions can also backfire by drawing attention to LLM presence and researcher scrutiny. In doing so, they may heighten participants’ self-consciousness or suspicions about other participants’ LLM usage, fueling the very demand effects and signalling behaviours they aim to prevent. In this way, strategies designed to reduce LLM Mediation and LLM Delegation may inadvertently exacerbate the LLM Spillover loop. To counteract this risk, mitigation must go beyond technical safeguards and reinforce shared norms around fairness, effort, and genuine participation. Establishing these norms may be crucial for re-anchoring expectations and sustaining the long-term integrity of online research.
@@
As the burden on individual researchers to detect and prevent LLM Pollution continues to grow, platform- and community-level support will be essential to making mitigation scalable and sustainable. Platforms hosting online research should take greater responsibility for ensuring data integrity. For example, by strengthening terms of service to prohibit unauthorised LLM use and by providing clearer participant guidance to align expectations. Features such as refund policies or abuse reporting tools may further incentivise responsible behaviour and motivate institutional investment in prevention. Beyond platform measures, fostering community-wide standards and practices is also important. Sharing knowledge, coordinating responses, and working toward common safeguards can reduce implementation burdens and improve consistency across the community. In the long term, reinvesting in physical lab infrastructure or supervised environments may be necessary in cases where higher control is needed. Such environments offer a level of oversight over participants’ devices, attention, and behaviour that remains difficult to achieve through technical safeguards alone. However, given the costs and practical limitations of lab-based research, collaborations will again be key to ensuring feasibility and broader access.

### v2

@@ Abstract
Online behavioural research faces an emerging threat as participants increasingly turn to large language models (LLMs) for advice, translation, or task delegation: LLM Pollution. This Perspective identifies three interacting variants through which LLM Pollution jeopardises the validity and integrity of online behavioural research. First, Partial LLM Mediation occurs when participants make selective use of LLMs for specific aspects of a task, such as translation or wording support, leading researchers to (mis)interpret LLM-shaped outputs as human ones. Second, Full LLM Delegation arises when agentic LLMs complete studies with little to no human oversight, undermining the central premise of human-subject research at a more foundational level. Third, LLM Spillover signifies human participants altering their behaviour as they begin to anticipate LLM presence in online studies, even when none are involved. While Partial Mediation and Full Delegation form a continuum of increasing automation, LLM Spillover reflects second-order reactivity effects. Together, these variants interact and generate cascading distortions that compromise sample authenticity, introduce biases that are difficult to detect post hoc, and ultimately undermine the epistemic grounding of online research on human cognition and behaviour. Crucially, the threat of LLM Pollution is already co-evolving with advances in generative AI, creating an escalating methodological arms race. To address this, we propose a multi-layered response spanning researcher practices, platform accountability, and community efforts. As the challenge evolves, coordinated adaptation will be essential to safeguard methodological integrity and preserve the validity of online behavioural research.
Raluca Rilla1†, Tobias Werner1,2†, Hiromu Yakura1†, Iyad Rahwan1, Anne-Marie Nussberger1†∗
@@
2University of Southampton, Department of Economics, Southampton SO17 1BJ, United Kingdom
@@
Fast, scalable access to diverse participants through online recruitment platforms such as Prolific or MTurk has enabled a revolution in behavioural research (Gosling and Mason, 2015; Goodman and Wright, 2022; Palan and Schitter, 2018). The promise of “100% human” responses lies at the core of their appeal for researchers seeking to study authentic human cognition and behaviour. [^1] For instance, Prolific.com, one of the leading survey providers, advertises its services as “Easily examine human actions and behaviours with our 100% human, ID-checked participants” (see https://www.prolific.com/academic-researchers, last accessed July 13th, 2025. Yet as large language models (LLMs) become increasingly capable and readily available, this promise is under growing threat, evidenced by participants themselves reporting routinely using LLMs to assist with study tasks (Zhang et al., 2025). We term this emerging phenomenon, in which LLMs are involved in online tasks intended to measure human responses, LLM Pollution. It represents a cause of several downstream effects on the integrity of online behavioural research that we discuss in this paper.
Recent experiences from our lab illustrate the immediacy of this threat: we repeatedly observed evidently LLM-mediated responses in up to 45% of submissions, characterised by overly verbose summaries of instructions or distinctly non-human phrases such as “I don’t experience confusion in the same way humans do” in open-ended survey questions. These observations are also corroborated by recent work from Veselovsky and colleagues (2025), who found that even when participants were explicitly asked not to use LLMs in a crowdsource task, a substantial share (up to 24%) still did so. Moreover, automated classifiers, behavioural heuristics, and self-reports indicated differing prevalence estimates, emphasizing how difficult it is to detect and prevent LLM Pollution reliably.
While threats to data quality from automated bots are not entirely new, earlier concerns focused on technically less sophisticated and more easily detectable bots (Goodrich et al., 2023; Griffin et al., 2022; Pozzar et al., 2020; Storozuk et al., 2020; Xu et al., 2022; Cresci, 2020). LLMs’ growing fluency and accessibility have, however, substantially amplified the problem by generating responses that are increasingly indistinguishable from those written by humans (Liem, 2025). As such, LLM Pollution can arise through a range of participant behaviours (Zhang et al., 2025). Some may use LLMs innocently, for example, to translate instructions or to improve fluency. Others might rely on them more strategically, using them to automate responses or reduce effort on tedious tasks. In more extreme cases, participants may delegate the completion of entire studies to agentic LLMs, such as OpenAI’s ChatGPT Agent, or open-source tools like Browser-use that can operate browsers, interpret screenshots, and navigate experiments without human oversight.
@@
Box 1 — A Case Study in LLM Pollution Detection and Incidence In a recent study we conducted with participants recruited via Prolific, we found that a lack of bot protection measures led to an alarming rate of seemingly bot-generated answers. This example is used here illustratively to demonstrate the types of challenges that can arise in practice rather than to report new empirical findings. In our initial pilot study, we only tracked copying and pasting on a page containing an open-ended question. Notably, 45% of participants engaged in one or both of these actions, suggesting potential cases of Partial LLM Mediation and Full LLM Delegation. Even to an untrained eye, some of these responses were obviously generated by LLMs, representing, for instance, cases of overly-abstract summarisations of all previous instructions. Given these results, we implemented the following measures for the main study: Implemented two reCAPTCHA v2 checkpoints at the beginning and middle of our survey, asking participants to confirm that they are not bots by checking a box and potentially completing a short validation test (0.2% failed to advance); Activated Qualtrics’ version of reCAPTCHA v3, which uses information about users’ behaviour and past activity to assign each a score between 0 and 1, with lower scores indicating likely bot activity (2.7% participants scored lower than 0.7, a threshold we determined on the basis of our piloting, with the lowest score being 0.2); Added a honeypot question that included white text invisible to humans, but readable to bots scraping all content (1.6% participants included our study-irrelevant keyword “hazelnut”; out of these 16 participants, only 2 had low reCAPTCHA v3 scores); Prevented copy-pasting, while still tracking which participants attempted these actions (4.7% participants attempted to copy and/or paste text). While these incidence rates were less concerning than those observed during piloting, we still identified identical nonsensical responses from a few different participants who otherwise appeared legitimate, having passed our honeypot question, having obtained perfect reCAPTCHA scores, and not having been flagged for copying/pasting. This suggests that even the above measures were insufficient for detecting different forms of LLM Pollution, further emphasising the scope of the problem.
In the following, we identify three key variants through which LLM Pollution manifests: Partial LLM Mediation, Full LLM Delegation, and LLM Spillover. We then outline a set of mitigation strategies, ranging from practical design choices to broader institutional safeguards. Finally, we offer an outlook on the future of online experiments in an era where human and machine responses are increasingly entangled. By synthesizing early anecdotal evidence and conceptual distinctions emerging from our work, this Perspective provides a forward-looking framework to guide cumulative empirical testing and collaborative, iterative refinement of mitigation strategies as LLM Pollution evolves.
@@
Figure 1: Three Variants of LLM Pollution. Partial LLM Mediation refers to cases where participants use LLMs for translation, idea generation, or performance gains. Full LLM Delegation involves the use of agentic tools or plugins to automate participation entirely. LLM Spillover refers to normative or behavioural shifts prompted by participants’ beliefs about LLM involvement, regardless of whether any is present. These variants can interact and reinforce one another, compounding their threat to research validity and inference. Illustration by H. Jahani.
@@
First, Partial LLM Mediation arises when participants rely on LLMs to help refine or generate their responses. Such responses may then appear to be human-generated but are, in part, machine-shaped. For instance, participants may rely on LLMs to translate content, improve the fluency of their writing, reduce cognitive effort, provide strategic advice on how to complete the task efficiently, or to maximise payment. Participants may copy and paste study instructions into LLM interfaces such as ChatGPT, or use screen-reading agents like Gemini Live to receive guidance and mediate their participation in the experiment. In some open-ended responses, LLM involvement may be evident from overly fluent language or stylistic quirks typical of LLMs (see, for instance, Yakura et al., 2024; Liang et al., 2024). Still, in many cases, Partial LLM Mediation remains difficult to detect from output alone, particularly when assistance is used only sparsely or intermittently. Partial LLM Mediation can compromise inference in several ways: because LLM-generated responses tend to exhibit lower variance and reflect dominant patterns in training data, their inclusion may artificially inflate central tendencies and obscure the true distribution of human responses (Zhang et al., 2025). These dominant patterns often encode systematic biases, such as the overrepresentation of Western, Educated, Industrialised, Rich, and Democratic (WEIRD) linguistic and cultural norms (Atari et al., 2023; Tao et al., 2024), which in turn are posed to produce homogenised outputs that mask meaningful variation, even within diverse participant samples. The decline of this naturally occurring diversity compromises the ecological and external validity of research findings, as responses reflect the linguistic regularities of training data rather than those of actual human populations. Researchers may believe they are analysing purely human-generated data, when in fact they are observing a blend of human and AI responses. This leads to incorrect inferences when the goal is to understand human cognition or behaviour. At the same time, Partial LLM Mediation has the potential to decrease internal validity by changing the very cognitive and linguistic processes that experimental manipulations aim to measure. Attempts to contrast human behaviour with machine behaviour may therefore rely on data that no longer represents the former in a clear manner.
@@
The second variant, Full LLM Delegation, involves participants outsourcing the entire study interaction to LLM-based tools or agents. These systems combine a language model with a control interface that enables goal-directed interaction with web environments. In practice, browser-integrated LLM agents (e.g., OpenAI’s ChatGPT Agent, open-source solutions like Browser-Use, Skyvern, or Nanobrowser) can autonomously read study content, interpret instructions, click through consent forms, complete questionnaires, and generate both structured and open-ended responses — all with minimal or no human intervention once launched. This form of delegation is made even easier by new, dedicated AI browsers like Perplexity’s Comet and ChatGPT Atlas, which build agentic functionality directly into the browser. Moreover, these agentic LLMs are increasingly capable of identifying instructions that are designed to mislead or manipulate their behaviour (see Figure B3). They are built to operate autonomously but are also designed to permit temporary human handover when they encounter barriers, such as CAPTCHA or other bot-detection systems. Furthermore, participants often have the option to instruct LLMs to adopt specific standpoints, personas, or tones that make their responses appear more credible both in terms of content and formatting at the surface level (Höhne et al., 2025a), further obscuring their underlying homogeneity (Zhang et al., 2025). Full LLM Delegation amplifies several risks observed in Partial LLM Mediation, as it more fully caters to participants’ motivations for efficiency and convenience, while distancing the resulting responses even further from genuine human behaviour. Hence, it first and foremost deepens the epistemic threat: researchers can no longer assume responses reflect human cognition, even probabilistically. Second, Full LLM Delegation may undermine treatment orthogonality, as LLMs may systematically respond differently across experimental conditions — especially in novel tasks — based on learned heuristics from training data (Mei et al., 2024; Schmidt et al., 2024). Third, Full LLM Delegation enables scale: participants (or third parties) can deploy LLM agents across multiple studies simultaneously, transforming data collection platforms into sites of automated labour arbitrage. As these systems continue to evolve and adapt rapidly, Full LLM Delegation is likely to become more common and harder to detect, posing a growing challenge for future research. Though described separately for clarity, Partial LLM Mediation and Full LLM Delegation lie on a continuum of increasing automation, with gradual shifts and no clear cutoffs between levels of LLM involvement.
The third variant, LLM Spillover, considers second-order effects of LLMs on human participants. Unlike Partial LLM Mediation or Full LLM Delegation, which focus on cases where LLMs are directly involved, LLM Spillover acknowledges that people’s mere expectations, prior experiences, and ambient awareness of LLMs may also shape their behaviour, even in the absence of direct model interactions. As such, LLM Spillover is likely to manifest in various ways; here, we focus on three imminent manifestations to illustrate its scope: Anticipatory Spillover occurs when participants expect LLMs to be present and adjust their behaviour accordingly. For instance, they might suspect that an interaction partner in a strategic game uses an LLM, or that the experimenter paired them with an LLM “partner” rather than a human one, and in response change their behaviour in terms of cooperation, trust, or disclosure. Residual Spillover can arise from prior exposure to LLMs in unrelated contexts. Such cross-domain effects can result from well-known mechanisms like linguistic or emotional priming and fixation, which have also been observed in other digital settings. For example, exposure to AI-suggested wording can shift people’s tone and language use beyond the moment of interaction (Hohenstein et al., 2023), and brief LLM exposure has been shown to reduce idea diversity and hinder later unassisted performance (Kumar et al., 2025). Similar carryover patterns have also been found in digital search and ideation tasks that narrow cognitive focus (Oppenheimer and Patterson, 2025). Within the context of Residual Spillover from LLMs, recent empirical work indicates that people’s spoken language has already shifted since the rollout of ChatGPT (Yakura et al., 2024), suggesting that we should expect both altered linguistic patterns in participant responses as well as signalling behaviours that seek to avoid overtly ‘LLM-like’ phrasing. Meanwhile, Normative Spillover reflects broader cultural adjustments in response to the rising ubiquity of LLMs, which might plausibly lead participants to perceive their use as (increasingly) acceptable.
Together, the three variants of LLM Pollution capture interrelated processes that are better understood in relation than isolation. Mitigation strategies should reflect their interdependence, as interventions aimed at one variant can inadvertently amplify another. For instance, tightening controls to curb LLM Delegation may heighten anticipatory spillover effects. Addressing LLM Pollution thus requires strategies that are effective, proportionate, and sensitive to participant experience. Ideally, such strategies will also be scalable, empirically testable, and resilient to the continued evolution of the threat. In the following, we share examples of measures we have found promising and which we hope can provide a basis for a continuous community effort to mitigate LLM Pollution.
@@
Since efforts to detect LLM Pollution are already headed towards an arms race — as tools improve, so too do evasion strategies — eliminating LLM-generated responses entirely is unlikely to be feasible. Instead, we propose pragmatic, multi-layered mitigation strategies aimed at raising the cost and lowering the feasibility of LLM Pollution across different levels of behaviour research ecosystems. This includes safeguards that individual researchers can implement in study design, greater responsibility on research platforms to ensure sample integrity, and broader efforts to foster community-wide norms that promote data quality. Crucially, different forms of LLM Pollution require tailored responses, as no single tool addresses all threats equally (Höhne et al., 2025b; Veselovsky et al., 2025).
@@
Measure Example Implementations Responsibility Purpose Variants Risks & Concerns
@@
Human-supervised or multimodal response collection Block copying and pasting, require spoken or webcam-enabled responses, or conduct live sessions where participants provide verbal answers (e.g., via Zoom) Researchers Prevention LLM Mediation LLM Delegation LLM Spillover Audio and video recording might have privacy concerns despite its effectiveness.
@@
Honeypot questions Add very small or white text with specific instructions for LLMs; implement prompt injection techniques Researchers Post-hoc detection LLM Mediation LLM Delegation Advanced LLM-based agents might be able to detect and ignore honeypot questions.
@@
One set of strategies focuses on preventative measures designed to increase friction for LLM use. Commercial bot protection tools like reCAPTCHA or Cloudflare are widely available and relatively easy to implement. These tools can block access from known LLM platforms or agents (e.g., OpenAI’s ChatGPT Agent), and the platforms themselves are less likely to override such protections for ethical reasons. However, locally run tools or hybrid user-agent models can often bypass these barriers by temporarily returning control to a human. Another line of defence is to vary the mode of instruction presentation by using screenshots, images, or short videos to reduce the ease of copying and pasting into LLMs. While this may limit casual use, vision-enabled models and screen-sharing tools can again circumvent such barriers. Relatedly, restricting input methods can make automation more difficult by means of disabling pasting, requiring typed responses, or blocking right-click functionality. Recent evaluations of explicit requests not to use LLMs and interface-based copy-paste restrictions indicate that these measures can reduce LLM-usage at least to some extent (Veselovsky et al., 2025). Also, using multimodal channels (e.g., collecting responses via voice input or recording video via webcam) may offer an additional layer of friction, as it is currently harder to automate reliably. In some cases, researchers may also opt to collect responses in real time through supervised sessions via Zoom, for example, where participants provide verbal answers that are directly recorded by experimenters. While this approach offers the highest level of verification, it substantially increases the human resources demands for labs. Importantly, all of these strategies must be weighed against usability and privacy concerns, particularly in longer studies or with vulnerable populations. Finally, researchers may design tasks that contemporary LLMs struggle with, such as visual illusions or Theory of Mind questions, combined with additional comprehension checks (see Figures A1 and A2 for examples). While current models tend to overcorrect or produce obviously wrong answers in such tasks, these vulnerabilities are likely to diminish as models improve. As such, preventative strategies may offer only temporary protection unless regularly updated.
@@
Beyond prevention, researchers can also take steps to identify LLM involvement after the fact. Honeypot questions, such as inserting white or inaccessible text, can detect whether a participant’s browser is scraping the entire page, as many LLM agents do. These traps exploit the tendency of automated tools to process all on-screen content indiscriminately, including elements that human users cannot see (see Appendix A.7 for examples). They can be combined with specific prompt injection techniques, such as jailbreaking or prompt leaking (Höhne et al. 2025b; see Appendix A.7 for examples and discussion). Recording behavioural data like typing speed, mouse movement, and tab switching can also flag suspicious patterns. While helpful, these signals can be mimicked by more advanced agents. Researchers may choose to run AI-generated text detectors on open-ended responses. These tools, such as GPTZero or Undetectable AI, can provide estimates of LLM involvement, but they often lack transparency and may quickly become outdated as models evolve.
@@
Mitigating human reactivity effects that come to bear in LLM Spillover poses a distinct challenge, as interventions themselves can influence participant expectations. Simply informing participants that bot detection tools are in use may discourage casual or opportunistic reliance on LLMs, and explicit reminders that responses should reflect their own judgment can reduce mild or exploratory use. However, these interventions can also backfire by drawing attention to LLM presence and researcher scrutiny. In doing so, they may heighten participants’ self-consciousness or suspicions about other participants’ LLM usage, fuelling the very demand effects and signalling behaviours they aim to prevent. In this way, strategies designed to reduce LLM Mediation and LLM Delegation may inadvertently exacerbate the LLM Spillover loop. To counteract this risk, mitigation must go beyond technical safeguards and reinforce shared norms around fairness, effort, and genuine participation. Establishing these norms may be crucial for re-anchoring expectations and sustaining the long-term integrity of online research.
@@
As the burden on individual researchers to detect and prevent LLM Pollution continues to grow, platform- and community-level support will be essential to making mitigation scalable and sustainable (Hitches et al., 2024). Platforms hosting online research should take greater responsibility for ensuring data integrity. For example, by strengthening terms of service to prohibit unauthorised LLM use and by providing clearer participant guidance to align expectations. Features such as refund policies or abuse reporting tools may further incentivise responsible behaviour and motivate institutional investment in prevention. Beyond platform measures, fostering community-wide standards and practices is also important. Sharing knowledge, coordinating responses, and working toward common safeguards can reduce implementation burdens and improve consistency across the community. In the long term, reinvesting in physical lab infrastructure or supervised environments may be necessary in cases where higher control is needed. Such environments offer a level of oversight over participants’ devices, attention, and behaviour that remains difficult to achieve through technical safeguards alone. However, given the costs and practical limitations of lab-based research, collaborations will again be key to ensuring feasibility and broader access.